OpenStudy (anonymous):

145% of what number is 116?

OpenStudy (anonymous):

it is 168.2

OpenStudy (anonymous):

It is not 168.2,

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

that is the wrong answer

OpenStudy (anonymous):

Notice that 145% shows that the number is larger than the number we are looking for.

OpenStudy (anonymous):

168.2 is 145% of 116

OpenStudy (anonymous):

It has to be lower than 116.

OpenStudy (anonymous):

So, we're going to reverse the way you find a percentage of a number. Instead of multiplying, divide.

OpenStudy (anonymous):

is it 8%?

OpenStudy (anonymous):

It's 80, because 80*1.45=116; 116/1.45=80

OpenStudy (anonymous):

or 80%?

jigglypuff314 (jigglypuff314):

145% as a decimal is 145/100 = 1.45 1.45x = 116 x meaning the number you're looking for so x = 116 divided by 1.45 = ?

jigglypuff314 (jigglypuff314):

80 would be correct :) just not percent

OpenStudy (anonymous):

lol read your answer after I posted but atleast I can say I got it right on my own and wasnt just given the answer :p

OpenStudy (anonymous):

Congrats! :D